Simon drew a rectangle whose width is 6 inches less than its length. The perimeter of the rectangle is 6 feet. What are the dimensions of the rectangle Simon drew?I need length and width in inches! Thanks a Bunch!
Answer by stanbon(48545)

6 feet = 72 inches
---------
Simon drew a rectangle whose width is 6 inches less than its length. The perimeter of the rectangle is 6 feet. What are the dimensions of the rectangle Simon drew?I need length and width in inches
--------------------------
Width = L-6
-------------------
Equation:
Perimeter = 2(L + W)
72 = 2(L + L-6)
36 = 2L-6
2L = 42
L = 21 inches (length)
---
W = L-6 = 15 inches (width)
==================================
Cheers,
Stan H.